OpenStudy (anonymous):

A plane in a level flight at 600 m travels at 150 mps horizontally. Its target ship is sailing at 20 mps in the same direction and on the same vertical plane as the bombing plane. At what horizontal distance behind the ship should a bomb be released in order to score a direct hit?

OpenStudy (anonymous):

please help....

OpenStudy (anonymous):



OpenStudy (anonymous):

i have actually computed, but i dunno if it is teh correct answer....is it -863.46 m?

OpenStudy (anonymous):

Okay, the most easy way to do this is taking reference where v ship s zero. The plane velocity then is 130 m/s relative to the ship. Time needed for the bomb to fall is \[t=\sqrt{\frac{2h}{g}}\] and the distance is \[x=v_{ps}t\\x=130\sqrt{\frac{2\times 600}{9.8}} \]

OpenStudy (anonymous):

i am confused with how u got 130m/s? i got the time but what i did to get the velocity is: x (plane)=(150)(11.07)=1660.5 then since the ship is more distant from the reference point, what i did to get the displacement between plane and ship is: 1660.5+x=(20)(11.07)=-1439.1. i got this answer. isn't this correct?

OpenStudy (anonymous):

oh, the answer with you solution is almost the same to that of my solution....thanks for verification =))))